mhanifa:

Given two independent events: P(A) = 2/9 P(B) = 4/11 Can you check if the answers are correct. (a) P(A and B) (the answer is 8/99) (b) P(not A and not B) (the answer is 91/99) (c) P(at least one event happens) (the answer is 48/99)

HighIQReptilian:

These all seem good! When finding the probability of both happening you multiplied (a), when finding the probability of neither you subtracted 8 from the total chances to give you it (b), and lastly you added the two probabilities to give you (c).

mhanifa:

Thanks!

mhanifa:

I think the last one is incorrect.

surjithayer:

C. \[P(does nothappen)=P(notA)P(notB)=\frac{ 7 }{ 9 }\times \frac{ 7 }{ 11 }=\frac{ 49 }{ 99}\] \[P(at~least~one~event~happens)=1-\frac{ 49 }{ 99 }=?\] i may be incorrect.

mhanifa:

Thanks! It means (b) is incorrect as well?

surjithayer:

i think so.

mhanifa:

So (a) is correct: 8/99 (b) should be 49/99 (c) should be 50/99

surjithayer:

yes

mhanifa:

Thanks!

surjithayer:

yw